what is the equation of the line that contains the point (-1,-2) and is perpendicular to the line determined by 2x-3y=6?
Answer by mananth(16946)

2 x -3 y = 6
Find the slope of this line

-3 y = -2 x + 6
Divide by -3
y = 0.667 x + -2
Compare this equation with y=mx+b
slope m = 0.67

The slope of a line perpendicular to the above line will be the negative reciprocal -1.5
m1*m2=-1
The slope of the required line will be

m= -1 1/2 ,point ( -1 , -2 )
Find b by plugging the values of m & the point in
y=mx+b
-2 = 1.50 + b
b= -3.5
m= -3/2
The required equation is y=-3/ 2 x-7/ 2
m.ananth@hotmail.ca